Estimated activity times and precedences are given below:

Activity Optimistic Most Likely Pessimistic Required Precedence
A 6 7 14
B 8 10 12
C 2 3 4
D 6 7 8 A
E 5 5.5 9 B, C
F 5 7 9 B, C
G 4 6 8 D, E
H 2.7 3 3.5 F

What is the probability that the project will be completed within:

A. 21 Days

B. 22 Days

C. 25 Days
Answer #1

Solution:

Activity Optimistic time-a Expected completion time-m Pessimistic time-b Expected time= (a+4*m+ b)/6 Variance, (sigma)^2= (b-a/6)^2
A 6 7 14 8.00 1.78
B 8 10 12 10.00 0.44
C 2 3 4 3.00 0.11
D 6 7 8 7.00 0.11
E 5 5.5 9 6.00 0.44
F 5 7 9 7.00 0.44
G 4 6 8 6.00 0.44
H 2.7 3 3.5 3.03 0.02
Path Duration
ADG 21.00
BEG 23.00
BFH 20.03
CEG 15.00
CFH 13.03

Expected project duration is 23 days through BEG

Answer: the probability that the project will be completed within:

A. 21 Days: 0.0416

step 1 we will find the variance of the tasks which lie on critical path 1.33
step 2 mean project time (u) of critical path is= 23.00
step 3 Required completion time is 21
step 4 standard deviation= sqrt(variance)=   1.155
step 5 because Z= (given completion time- u)/standard deviation -1.732
step 6 P(z)= using NORM.S.DIST(z,true) 0.0416

Answer: the probability that the project will be completed within:

B. 22 Days: 0.1932

step 1 we will find the variance of the tasks which lie on critical path 1.33
step 2 mean project time (u) of critical path is= 23.00
step 3 Required completion time is 22
step 4 standard deviation= sqrt(variance)=   1.155
step 5 because Z= (given completion time- u)/standard deviation -0.866
step 6 P(z)= using NORM.S.DIST(z,true) 0.1932

Answer: the probability that the project will be completed within:

C. 25 Days: 0.9584

step 1 we will find the variance of the tasks which lie on critical path 1.33
step 2 mean project time (u) of critical path is= 23.00
step 3 Required completion time is 25
step 4 standard deviation= sqrt(variance)=   1.155
step 5 because Z= (given completion time- u)/standard deviation 1.732
step 6 P(z)= using NORM.S.DIST(z,true) 0.9584

Please ask, if you have any doubts through the comment section

Do rate the answer